OpenStudy (anonymous):

w=Cr^-2 Solve for r.

OpenStudy (anonymous):

r= sqr(C/w)

OpenStudy (anonymous):

\[w = Cr^{-2}\] or is it \[w = (Cr)^{-2}\] ill assume the first \[w = \frac{C}{r^2}\] \[r^2 = \frac{C}{w}\] \[r = \sqrt{\frac{C}{w}}\]

OpenStudy (anonymous):

it was the first there was no parenthesis

OpenStudy (anonymous):

can you tell me the steps you did?

OpenStudy (anonymous):

rewrote \[w = Cr^{-2}\]as\[w=\frac{C}{r^2}\]

OpenStudy (anonymous):

\[\text{multiplied by }r^2 \text{ and divided by }w \] \[r^2=\frac{C}{w}\] then I took the square root\[r = \sqrt{\frac{C}{w}}\]

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

b. Suppose that an object is 100 pounds when it is at sea level. Find the value of C that makes the equation true. (Sea level is 3,963 miles from the center of the Earth.)

OpenStudy (anonymous):

what would i do first?

OpenStudy (anonymous):

\[C=100x(3963x3963)\]